A sample of 0.536mol of Nitrogen oxide gas occupies a volume of 0.514L.What would be the volume if the number of moles of nitrogen oxide gas is increase to 0.883 with no change in temperature and pressure

a.1.00L

b.0.847L

c.0.229L

d.0.923L
Answer #1

by ideal gas equation

PV=nRT

when pressur and temperature are constant

V2/V1 = n2/n1

V/0.514 = 0.883/0.536

V=0.847L (A)